# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversational Agents

### v4

@@ Appendix I Experimental Results in an Unlimited Time Setting
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. Open Tülu 2-70B 2.17 (0.46) 27.24 (1.17) 33.34 (1.17) 32.85 (1.85) 34.95 (0.47) 29.41 (1.22) 33.55 (2.79) Tülu 2-7B 0.63 (0.26) 30.26 (1.03) 27.68 (1.24) 30.98 (1.08) 30.99 (0.22) 27.93 (1.97) 31.80 (2.05) Llama3.1-70B 31.03 (1.91) 28.91 (2.33) 38.44 (5.98) 41.68 (3.68) 38.40 (1.10) 40.83 (1.07) 44.27 (0.57) Llama3.1-8B 26.65 (1.19) 25.80 (0.18) 32.01 (1.10) 30.48 (1.30) 29.50 (1.10) 32.82 (0.63) 32.23 (2.32) Mixtral-8x7B 38.92 (1.61) 26.91 (1.46) 39.98 (2.98) 27.02 (0.42) 30.27 (1.37) 35.89 (0.37) 29.52 (1.28) Mistral-7B 29.10 (1.34) 23.33 (0.83) 34.59 (0.80) 27.87 (2.66) 30.59 (2.09) 30.45 (0.89) 27.93 (0.99) Gemma-7B 17.37 (0.77) 22.58 (1.62) 21.41 (1.53) 21.61 (1.53) 23.90 (0.90) 21.61 (1.09) 20.88 (0.91) Gemma-2B 0.37 (0.07) 23.48 (1.85) 23.13 (1.14) 25.72 (2.67) 29.12 (1.90) 24.88 (1.40) 24.81 (1.31)
$\dagger$: Llama3.1 supports up to 128k tokens, but we limited it to 8k tokens due to GPU VRAM limits.

### v5

@@ Appendix I Experimental Results in an Unlimited Time Setting
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 32.06 (0.52) 30.01 (1.09) 46.57 (1.56) 34.44 (1.53) 34.23 (0.99) 38.78 (0.24) 37.97 (1.24) Open Tülu 2-70B 2.17 (0.46) 27.24 (1.17) 33.34 (1.17) 32.85 (1.85) 34.95 (0.47) 29.41 (1.22) 33.55 (2.79) Tülu 2-7B 0.63 (0.26) 30.26 (1.03) 27.68 (1.24) 30.98 (1.08) 30.99 (0.22) 27.93 (1.97) 31.80 (2.05) Llama3.1-70B 31.03 (1.91) 28.91 (2.33) 38.44 (5.98) 41.68 (3.68) 38.40 (1.10) 40.83 (1.07) 44.27 (0.57) Llama3.1-8B 26.65 (1.19) 25.80 (0.18) 32.01 (1.10) 30.48 (1.30) 29.50 (1.10) 32.82 (0.63) 32.23 (2.32) Mixtral-8x7B 38.92 (1.61) 26.91 (1.46) 39.98 (2.98) 27.02 (0.42) 30.27 (1.37) 35.89 (0.37) 29.52 (1.28) Mistral-7B 29.10 (1.34) 23.33 (0.83) 34.59 (0.80) 27.87 (2.66) 30.59 (2.09) 30.45 (0.89) 27.93 (0.99) Gemma-7B 17.37 (0.77) 22.58 (1.62) 21.41 (1.53) 21.61 (1.53) 23.90 (0.90) 21.61 (1.09) 20.88 (0.91) Gemma-2B 0.37 (0.07) 23.48 (1.85) 23.13 (1.14) 25.72 (2.67) 29.12 (1.90) 24.88 (1.40) 24.81 (1.31)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.